OpenStudy (anonymous):

Factor x2 + 25x – 100. A. (x – 2)(x + 50) B. (x – 5)(x + 20) C. (x + 4)(x – 25) D. prime

OpenStudy (anonymous):

b

OpenStudy (acxbox22):

NO IT IS PRIME multiply b out and you get x^2+15x-100

OpenStudy (campbell_st):

lol.... well the problem would be that there is a possible type... or a correct answer isn't listed...

OpenStudy (john_es):

:)

OpenStudy (anonymous):

I put B first but it was prime

OpenStudy (campbell_st):

or... perhaps the quadratic is prime use the discriminant to check \[25^2 - 4 \times 1 \times 100\]

OpenStudy (campbell_st):

oops should read \[25^2 - 4 \times 1 \times -100 = \]

OpenStudy (anonymous):

Factor x2 + 6x + 8. A. (x + 1)(x + 6) B. (x + 1)(x + 8) C. (x + 2)(x + 4) D. prime

OpenStudy (anonymous):

i think C

OpenStudy (anonymous):

it is

OpenStudy (anonymous):

D prime